Two forces are acting on a boulder sitting on a fl
Answer #1

step:1

Given that

mass m=30 kg

force F1=12.2 N

force F2=14 N

step:2

now we find the net force in case 1

fig 1=>Fnet=[F1^2+F2^2]^1/2=[12.2^2+14^2]^1/2=18.6 N

acceleration a=Fnet/m=18.6/30=0.62 m/s^2

step:3

now we find the net force in case to

Fnet =[F2cos\theta]i+[F1+F2sin\theta] j

=[14 cos30]i+[12.2+14 sin30] j

=12.12 i+19.2 j

Fnet =[12.12^2+19.2^2]^1/2=22.71 N

acceleration a=22.71/30=0.76 m/s^2